OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7 HELP ME PLEASE

OpenStudy (triciaal):

let y = y then x + 7 = -2x + 19 solve for x substitute this value of x in either of the original to find y

OpenStudy (triciaal):

remember to check your answer x = 4 and y = 11

OpenStudy (idealist10):

-2x+19=x+7 19=x-(-2x)+7 19=x+2x+7 19=3x+7 3x=19-7 3x=12 x=12/3 x=4 Answer: 4

OpenStudy (triciaal):

good. now that you have x you need to find y

OpenStudy (triciaal):

solve the system means find both the x value and the y value that satisfies both equations

OpenStudy (idealist10):

Sorry, I forgot to solve for y. This is simple, just plug in x into any equation you want to solve for y. y=x+7 y=4+7=11 y=11 Answer: (4, 11)

OpenStudy (triciaal):

@Idealist10 I did not even check that you were not the one that posted. @kyshawn 123 I posted what to do and @Idealist10 did it for you.